### v1

# Wasserstein Stability of Persistence Landscapes and Barcodes

Abstract
Barcodes form a complete set of invariants for interval decomposable persistence modules and are an important summary in topological data analysis. The set of barcodes is equipped with a canonical one-parameter family of metrics, the $p$-Wasserstein distances. However, the $p$-Wasserstein distances depend on a choice of a metric on the set of interval modules, and there is no canonical choice. One convention is to use the length of the symmetric difference between 2 intervals, which equals to the 1-norm of the difference between their Hilbert functions. We propose a new metric for interval modules based on the rank invariant instead of the dimension invariant. Our metric is topologically equivalent to the metrics induced by the $p$-norms on $\mathbb{R}^{2}$. We establish stability results from filtered CW complexes to barcodes, as well as from barcodes to persistence landscapes. In particular, we show that vectorization via persistence landscapes is 1-Lipschitz with a sharp bound, with respect to the 1-Wasserstein distance on barcodes and the 1-norm on persistence landscapes.
1 Introduction
Topological Data Analysis (TDA) is a study of the shape of data using algebraic topology. Persistent homology is one of its main tools. A common setting involves data given in the form of a point cloud, from which one can construct a filtered chain complex. By taking homology one obtains a persistence module. Pointwise finite dimensional (p.f.d.) persistence modules can be decomposed into interval modules. This decomposition can be represented as a persistence diagram (PD) or a barcode, which encodes the birth and death of homological classes as the filtration parameter changes. It provides topological summaries used to interpret the shape of data.
The $p$-Wasserstein distances between persistence diagrams (PDs) depend on a choice of metric on the points in the PDs, which represent the interval modules in the interval decomposition. In Section 3, we introduce a new metric on the set of interval modules, denoted by $d_{\operatorname{rank}}$, which compares modules via their rank invariants. Standard choices of metrics for interval modules depend on the dimension invariant [6, 12, 33]. Our metric puts more weight on more persistent features by imposing a higher cost to move points further from the diagonal. The standard choices of metrics induce the same topology on $\mathbb{R}^{2}$ and are uniformly equivalent, whereas our metric $d_{\operatorname{rank}}$ is topologically equivalent while not uniformly equivalent to them. We will use the 1-Wasserstein distance together with $d_{\operatorname{rank}}$, which we denote $W_{1}^{\operatorname{rank}}$. Observe that for any metric $d$ and $p\leq q$, $W_{p}^{d}\geq W_{q}^{d}$, i.e. the 1-Wasserstein distance is the most discriminative distance among all $p$. We equip the set of PDs with $W_{1}^{\operatorname{rank}}$ and show stability from filtered chain complexes to PDs, as well as from PDs to persistence landscapes (PLs).
$\begin{split}\textbf{Filtered Chain Complexes}\;\longrightarrow&\;\;\textbf{Persistence Modules}\quad\longrightarrow\textbf{Persistence Landscapes}\\ \|\{d_{\operatorname{rank}}(I_{w(\sigma)},I_{v(\sigma)})\}_{\sigma\in K}\|_{1}\;\geq\;\;&W_{1}^{\operatorname{rank}}(\operatorname{Dgm(w)},\operatorname{Dgm(v)})\>\geq 2\|\Lambda(\operatorname{Dgm(w)})-\Lambda(\operatorname{Dgm(v)})\|_{1}\end{split}$
In Section 4, we prove a stability result from weighted chain complexes, or equivalently filtered chain complexes, to PDs. We choose an interval and view filtered chain complexes as persistence modules indexed by this interval. The interval is chosen to give the strongest stability result 4.13. We measure the distance between weights on a common chain complex using $d_{\operatorname{rank}}$.
In practice, one often needs to perform further analysis on topological summaries like PDs, using machine learning and statistical tools which take vectors as inputs. The space of PDs admits metric space structures but not vector space structure. Thus mapping into vector summaries is useful for such analyses. In Section 5, we consider the vectorization using persistence landscapes defined in [3] and show a sharp 1-Lipschitz stability result using the $W_{1}^{\operatorname{rank}}$ on PDs and $L^{1}$ norm on PLs.
In Section 6, we consider an alternative construction of PLs that applies the Möbius inversion of graded rank functions. This produces a sequence of signed PDs, called graded PDs [2]. We use the notion of graded Wasserstein distances to provide a stability result of persistence landscapes in Proposition 6.4.
1.1 Wasserstein distances and metric for interval modules
For $1\leq p\leq\infty$, the p-Wasserstein distances depend on a choice of metric on the points in the PDs, and there is no canonical choice. A PD is a representation of the interval decomposition of a persistence module, and each point in the PD represents an interval module in the decomposition. In particular, interval modules are objects in the functor category from the poset $\mathbb{R}$ to finite dimensional vector spaces, which we denote $\mathsf{vect}^{\mathbb{R}}$.
Let $M$ be a p.f.d persistence module. Let $\mathbb{R}_{\leq}^{2}=\{(x,y)\in\mathbb{R}^{2}:x\leq y\}$. The dimension and rank invariant of $M$ are functions given below.
$\begin{split}\operatorname{dim}M:\mathbb{R}&\to\mathbb{Z}\\ r&\mapsto\operatorname{dim}M_{r}\end{split}$
$\begin{split}\operatorname{rank}M:\mathbb{R}_{\leq}^{2}&\to\mathbb{Z}\\ (r,s)&\mapsto\operatorname{rank}M_{r\leq s}\end{split}$
Figure 1: Let $I\oplus J$ denote the direct sum of two interval modules for two overlapping intervals II and $J$. We plot the function $\operatorname{rank}\>I\oplus J:\mathbb{R}_{\leq}^{2}\to\mathbb{Z}$. By restricting the domain to the diagonal, the rank function recovers the Hilbert function, i.e. $\dim I\oplus J(t)=\operatorname{rank}I\oplus J(t,t)$.
To define a metric on interval modules, we compare their invariants via function norms. Below are some metrics used previously in the TDA community. In Section 3, we reformulate them from the perspective of invariants and function norms and introduce a new metric $d_{\operatorname{rank}}$.
  1. 1.
    $d_{\infty}(I,J)=\|(\inf I,\sup I)-(\inf J,\sup J)\|_{\infty}$.
    The bottleneck distance is defined using $d_{\infty}$. Some also use $d_{\infty}$ for $p$-Wasserstein distances [1, 30, 20].
  2. 2.
    $d_{\operatorname{dim}}(I,J)=\text{length}(I\Delta J)=\int_{\mathbb{R}}\operatorname{dim}I-\operatorname{dim}J|$ [6]
Definition 1.1.
Let $d_{\operatorname{rank}}(I,J)=\int_{\mathbb{R}_{\leq}^{2}}|\operatorname{rank}\>I-\operatorname{rank}\>J|$. Let $W_{1}^{\operatorname{rank}}$ denote the 1-Wasserstein distance with metric $d_{\operatorname{rank}}$.
Note the rank invariants are complete invariants in $\mathsf{vect}^{\mathbb{R}}$ whereas the dimension invariants are not.
Our new metric assigns higher cost to moving points with greater persistence, in contrast to prior metrics, which depend only on the relative distance between points, regardless of their location in the diagram.
Figure 2: We plot the boundaries of the balls of radius $r$ centered at $x$ for each of $d_{\operatorname{rank}}$ (red), $d_{\operatorname{dim}}$ (blue), and $d_{\infty}$ (green). The radius increases from left to right.
1.2 More about dimension invariants, rank invariants, and metrics
If we plot the dimension functions on $\mathbb{R}$, we obtain what’s called the Betti curve. Betti curves contain strictly less information than barcodes, and there are different barcodes with the same betti curve. In other words, the betti curve tells no information about the persistence of homology classes.
Given a filtered topological space $X:\mathbb{R}\to\mathsf{Top}$, one can define its persistent homology and betti curve.
$H_{j}X:\mathbb{R}\to\mathsf{Vect}$
$\beta_{j}X:\mathbb{R}\to\mathbb{Z}$
One can express the metric $d_{\operatorname{dim}}$ using the betti curve.
$d_{\operatorname{dim}}(H_{j}X,0)=\int\beta_{j}X$
And $d_{\operatorname{dim}}(H_{j}X,0)=d_{\operatorname{dim}}(H_{j}Y,0)$ even if they have vastly different PDs.
1.3 Related work
1.3.1 Related work on barcode stability
There are a number of stability results for PDs based on different constructions. Cohen-Steiner et al. [13] studied PDs arising from Lipschitz functions on compact metric spaces and proved bottleneck stability. Gulen, McCleary, and Patel [17, 23] investigated stability with respect to the edit distance; in particular, Gulen and Patel [17] considered weighted PDs obstained from metric spaces equipped with fully supported probability measures. Both works employed functorial constructions of PDs and used functoriality to establish stability. Other results on stability from filtration functions to PDs include the work of Skraba et al. [33], which generalizes the bottleneck stability to p-Wasserstein stability by choosing $\ell^{p}$ as the ground metric for $W_{p}$. In section 4.1, we rephrase this stability result expressed explicitly by dimension invariants.
Let $f:\sigma\mapsto d_{\operatorname{dim}}(I_{w(\sigma)},I_{v(\sigma)})$ where $I_{w(\sigma)},I_{v(\sigma)}$ are free modules generated by $\sigma$ with respect to $w,v$. Observe f=wvf=w-v. We can rewrite the bottleneck stability[12] as:
$W_{\infty}^{d_{\infty}}(\operatorname{Dgm}(w),\operatorname{Dgm}(v))\leq\|f\|_{\infty}$
The $p$-Wasserstein stability of barcodes in [33] can be reformulated as:
$W_{p}^{d_{p}}(\operatorname{Dgm}(w),\operatorname{Dgm}(v))\leq\|f\|_{p}$
Let $g:\sigma\mapsto d_{\operatorname{rank}}(I_{w(\sigma)},I_{v(\sigma)})$. We rewrite theorem 4.8 as:
$W_{1}^{\operatorname{rank}}(\operatorname{Dgm}(w),\operatorname{Dgm}(v))\leq\|g\|_{1}$
1.3.2 Related work on vectorization stability
Bubenik proved stability results for PLs [3], using the bottleneck distance on PDs and $L^{\infty}$ norm on PLs. We provide a sharp 1-Lipschitz stability result using $W_{1}^{\operatorname{rank}}$ and the $L^{1}$ norm instead. Various vectorizations of barcodes have been proposed and proven stable [28, 32, 1, 7, 20]. Persistence images, Persistence scale space kernel, tropical coordinates [1, 30, 20] are shown to be Lipschitz w.r.t. $W_{1}$, all with ground metric $d_{\infty}$, and the Lipschitz constants depend on external parameters such as kernels, weights, scales.
1.3.3 Our contribution
We define a new metric on interval modules and show the PDs and PLs are $W_{1}$-stable. We prove a sharp 1-Wasserstein stability of PLs with the Lipschitz constant $\frac{1}{2}$. Unlike most other stability results for vectorizations, this Lipschitz constant does not depend on any external parameters.
2 Background
2.1 Notation
In this paper, $(P,\leq)$ is a finite totally ordered set. We also consider $(P,\leq)$ to be a category, where the objects are elements in $P$, and there exists a unique morphism from $p$ to $q$ whenever $p\leq q$. Let $\mathsf{Ch}_{\geq}$ denote the category of non-negatively indexed chain complexes, with chain maps as morphisms. Let $\mathbb{R}_{\leq}^{2}$ denote $\{(x,y)\in\mathbb{R}^{2}:x\leq y\}$ and $\Delta$ denote $\{(x,y):x=y\}$. Consider $\mathbb{R}_{\leq}^{2}$ with the subspace topology from the standard topology on $\mathbb{R}^{2}$. Equip $\mathbb{R}_{\leq}^{2}/\Delta$ with the quotient topology. Let $(\mathbb{R}^{2}_{\leq},\Delta)$ denote the pointed topological space $(\mathbb{R}^{2}_{\leq}/\Delta,\Delta)$.
Let $\operatorname{Int}(\mathbb{R})$ be the set of intervals in $\mathbb{R}$. Let $\mathbb{k}$ be a field. $\mathsf{Vect}_{\mathbb{k}}$ denotes the category of $\mathbb{k}$-vector spaces with $\mathbb{k}$-linear maps as morphisms, and we drop the field when clear in context. $\mathsf{Vect}^{\mathbb{R}}$ denotes the category consisting of functors from the poset category $(\mathbb{R},\leq)$ to $\mathsf{Vect}$.
2.2 Persistent homology
A filtered chain complex is a functor $F:\mathbb{R}\to\mathsf{Ch}_{\geq}$ such that $\forall a\leq b\in\mathbb{R},\mathsf{Ch}_{\geq}(a\leq b)$ is a collection of inclusion maps. Let $K$ be a finite cell complex. A weight on $K$ is a functor $w:(K,\subset)\to(\mathbb{R},\leq)$. We say $M$ is a persistence module if $M:\mathbb{R}\to\mathsf{Vect}$ is a functor. In other words, $M$ consists of a collection of vector spaces and linear maps compatible with compositions, $\{M_{r}:r\in\mathbb{R}\}$ and $\{M_{r\leq s}:r,s\in\mathbb{R}\}$. One common construction of a persistence module is to take homology of a filtered chain complex, $H_{i}(F-;\mathbb{k}):(\mathbb{R},\leq)\to\mathsf{Ch}_{\geq}\to\mathsf{Vect}_{\mathbb{k}}$.
Let $I\in\operatorname{Int}(\mathbb{R})$. The corresponding interval module $I$ is an object in $\mathsf{Vect}_{\mathbb{k}}^{\mathbb{R}}$, where $I_{r}=\mathbb{k}$ for $r$ in the interval $I$ and 0 otherwise, and $I_{r\leq s}=1_{\mathbb{k}}$ if $r\leq s$ and $r,s\in I$, and 0 otherwise. We say a persistence module $M$ is interval decomposable if $M\cong\underset{j}{\oplus}I_{j}$ where each $I_{j}$ is an interval module. Interval modules are indecomposable, meaning there is no isomorphism from an interval module $I$ to $\underset{j}{\oplus}N_{j}$ where $N_{j}\in\mathsf{Vect}^{\mathbb{R}}$. Interval decompositions are unique up to re-indexing.
Theorem 2.1.
(Structure Theorem in TDA [14]): Let $M$ be a persistence module. If $M$ is pointwise finite dimensional (p.f.d.), then $M$ is interval decomposable, and the interval decomposition is unique up to isomorphism.
If a persistence module $M$ is interval decomposable, then we can represent the interval decomposition using the following 2 methods:
  1. 1.
    barcodes: a multiset of intervals in $\mathbb{R}$ where each interval/bar is the support of an interval module in the decomposition.
  2. 2.
    persistence diagrams: a persistence diagram (PD) can be viewed as a formal sum $\alpha=\underset{j}{\sum}x_{j}$ for $x_{j}\in(\mathbb{R}_{\leq}^{2},\Delta)$ and $x_{j}=(\inf I_{j},\sup I_{j})$. Equivalently, a PD can be expressed as a multiplicity function $\operatorname{Dgm}(F):(\mathbb{R}^{2}_{\leq}/\Delta,\Delta)\to\mathbb{Z}$. The 2 expressions are useful in different contexts.
Definition 2.2.
Let $\alpha,\beta$ be PDs and $\alpha=\underset{i\in I}{\sum}x_{i}$, $\beta=\underset{j\in J}{\sum}y_{j}$. A coupling $\gamma$ between $\alpha,\beta$ is a partial matching between the points in $\alpha,\beta$ given as follows:
$\gamma=\underset{i\in I^{\prime}}{\sum}(x_{i},y_{\phi(i)})+\underset{i\in I\setminus I^{\prime}}{\sum}(x_{i},\Delta)+\underset{j\in J\setminus\phi(I^{\prime})}{\sum}(\Delta,y_{j})$
where $I^{\prime}\subset I,\phi:I^{\prime}\hookrightarrow J$.
Definition 2.3.
Let $d$ be a metric, $\gamma$ be a coupling. Define $\operatorname{Cost}(\gamma)=\int d(x,y)\>d\gamma$. When $\gamma$ is a coupling between discrete measures, such as PDs, $\gamma$ is a matching between points in the PDs with points on the diagonal. Then $\operatorname{Cost}(\gamma)=\underset{(x.y)\in\gamma}{\sum}d(x,y)$. More generally, define $\operatorname{Cost}_{p}(\gamma)=(\int d(x,y)^{p}\>d\gamma)^{1/p}$ for $p\in[1,\infty]$.
Definition 2.4.
The $p$-Wasserstein distance between PDs is given by: $W_{p}^{d}(\alpha,\beta)=\underset{\gamma}{inf}\>\operatorname{Cost}_{p}(\gamma)$.
Recall that $W_{p}^{d}$ satisfies the triangle inequality for finite PDs. In section 5, we extend the result to countable PDs.
2.3 Wasserstein distances for graded persistence diagrams
In the appendix, we recall the definition of Möbius inversion and the construction of graded PDs from [2]. This section contains the background knowledge for an alternative stability result using graded PDs in section 6.
Definition 2.5.
Let $\alpha,\beta:\mathbb{R}_{\leq}^{2}\to\mathbb{Z}$. Define the 1-Wasserstein distance to be $W_{1}^{d}(\alpha,\beta)=W_{1}^{d}(\alpha_{+}+\beta_{-},\beta_{+}+\alpha_{-})$, where $\alpha_{+}=\frac{|\alpha|+\alpha}{2},\alpha_{-}=\frac{|\alpha|-\alpha}{2},\beta_{+}=\frac{|\beta|+\beta}{2},\beta_{-}=\frac{|\beta|-\beta}{2}$.
Lemma 2.6.
Let $\alpha,\beta,\sigma,\tau$ be finite PDs. Let $d$ be a metric. Then
$W_{1}^{d}(\alpha+\sigma,\beta+\tau)\leq W_{1}^{d}(\alpha,\beta)+W_{1}^{d}(\sigma,\tau)$ (2.1)
and
$W_{1}^{d}(\alpha+\sigma,\beta+\sigma)=W_{1}^{d}(\alpha,\beta)$ (2.2)
If needed, please see proofs to equation 2.1 and 2.2 in appendix.
Corollary 2.7.
Let $\alpha,\beta,\sigma,\tau,\gamma,\eta$ be finite PDs. Let $d$ be a metric.
$W_{1}^{d}(\alpha+\gamma,\beta+\sigma)+W_{1}^{d}(\tau+\sigma,\eta+\gamma)\geq W_{1}^{d}(\alpha+\tau,\beta+\eta)$
The proof follows from 2.1 and $2.$
3 Wasserstein distances and metrics on interval modules
3.1 New metric on interval modules
The p-Wasserstein distances depend on a choice of metric on the points in the PDs, and there is no canonical choice. Conventional choices include metrics induced by the $p$-norm on $\mathbb{R}^{2}$, which can be interpreted using the dimension functions. In section 1.1, we introduced notations for 2 common metrics, $d_{\operatorname{dim}}$ and $d_{\infty}$, and express $d_{\operatorname{dim}}$ using the dimension function. We rewrite $d_{\infty}$ in terms of the dimension function in 4.10 later.
In fact, each invariant in $\mathsf{vect}^{\mathbb{R}}$ together with a functional norm defines a metric on the set of interval modules. For example, rank functions and Hilbert functions are in the normed function spaces equipped with $L^{p}$ norms for $1\leq p\leq\infty$. Each $L^{p}$ norm induces a metric on these functions of persistence modules.
Definition 3.1.
Let $d_{\operatorname{rank}}(I,J)=\int_{\mathbb{R}^{2}}|\operatorname{rank}\>I-\operatorname{rank}\>J|$. Let $W_{1}^{\operatorname{rank}}$ denote the 1-Wasserstein distance with metric $d_{\operatorname{rank}}$.
Lemma 3.2.
For interval modules $I,J$, let $T_{I}=\{(a,b)\in\mathbb{R}_{\leq}^{2}:\inf I\leq a\leq b\leq\sup I\}$ and $T_{J}$ likewise. Then, $d_{\operatorname{rank}}(I,J)=\text{Area}(T_{I}\Delta T_{J})$
Proof.
The proof follows from $|\operatorname{rank}I-\operatorname{rank}J|$ has value 1 on $T_{I}\Delta T_{J}$ and 0 otherwise. ∎
Our new metric assigns higher cost to moving points with greater persistence, in contrast to prior metrics, which depend only on the relative distance between points, regardless of their location in the diagram.
Observe that $d_{\operatorname{dim}}$ is more discriminative than $d_{\infty}$, but both are incomparable to $d_{\operatorname{rank}}$. And $W_{1}^{\operatorname{rank}}$ is incomparable to $W_{p}^{q}$ for any $p,q\in[1,\infty]$.
Lemma 3.3.
$d_{\operatorname{dim}}$ and $d_{\infty}$ are incomparable to $d_{\operatorname{rank}}$.
Figure 3: We plot the balls of radius $r$ centered at $x$ for each of $d_{\operatorname{rank}}$ (red), $d_{\operatorname{dim}}$ (blue), and $d_{\infty}$ (green). The radius increases from left to right.
As shown in figure 3, $d_{\operatorname{rank}}$ is incomparable to the other 2 metrics. Let $x\in(\mathbb{R}_{\leq}^{2},\Delta)$, representing the interval module $I_{x}$. The boundary $\{y:d_{\operatorname{rank}}(x,y)=r\}$ is given by 2 lines, each with constant persistence, and 2 hyperbolas. The 4 equations correspond to the following 4 cases.
Figure 4: Each ball is divided into 4 regions corresponding to the 4 cases. As the radius $r$ changes, each of the 4 curves can change convexity.
Let $x,y\in\mathbb{R}_{\leq}^{2}$ with coordinates $(x_{1},x_{2}),(y_{1},y_{2})$, corresponding to interval modules $I_{x},I_{y}$.
  1. 1.
    $I_{x}\subset I_{y}$
    $\frac{1}{2}(y_{2}-y_{1})^{2}-\frac{1}{2}(x_{2}-x_{1})^{2}=r$ (3.1)
  2. 2.
    $I_{y}\subset I_{x}$
    $\frac{1}{2}(y_{2}-y_{1})^{2}-\frac{1}{2}(y_{2}-y_{1})^{2}=r$ (3.2)
  3. 3.
    $I_{y}$ is to the right of $I_{x}$, i.e. there’s a nontrivial map $I_{x}\to I_{y}$
    $\frac{1}{2}(x_{2}-x_{1})^{2}+\frac{1}{2}(y_{2}-y_{1})^{2}-(x_{2}-y_{1})^{2}=r$ (3.3)
  4. 4.
    $I_{y}$ is to the left of $I_{x}$, i.e. there’s a nontrivial map $I_{y}\to I_{x}$.
    $\frac{1}{2}(x_{2}-x_{1})^{2}+\frac{1}{2}(y_{2}-y_{1})^{2}-(y_{2}-x_{1})^{2}=r$ (3.4)
Figure 5: The ball is defined by 4 curves, corresponding to the 4 cases. In particular when an interval doesn’t contain or being contained by $I_{x}$, the boundary curves are hyperbolas.
More generally, we can express $d_{\operatorname{rank}}(x,y)$ using lemma 3.2 and the Principle of Inclusion and Exclusion.
$d_{\operatorname{rank}}(x,y)=\frac{1}{2}(x_{2}-x_{1})^{2}+\frac{1}{2}(y_{2}-y_{1})^{2}-\max(\min(x_{2},y_{2})-\max(x_{1},y_{1}),0)^{2}$ (3.5)
3.2 Topological equivalence of metrics
We recall the definitions for topological equivalence and bilipschitz/uniformly equivalence between metrics. Let $d,d^{\prime}$ be 2 metrics on a set $X$. $d,d^{\prime}$ are topologically equivalent if $Id:(X,d)\to(X,d^{\prime})$ is a homeomorphism. $d,d^{\prime}$ are bilipschitz equivalent if $Id$ and $Id^{-1}$ are Lipschitz.
It is known that the metrics on $\mathbb{R}^{2}$ induced by the $p$-norms are bilipschitz equivalent. We prove the topological equivalence of $d_{\operatorname{rank}},d_{\operatorname{dim}}$ on $(\mathbb{R}_{\leq}^{2},\Delta)$ and disprove their bilipschitz equivalence.
Proposition 3.4.
$Id:(\mathbb{R}_{\leq}^{2}/\Delta,d_{\operatorname{rank}},\Delta)\to(\mathbb{R}_{\leq}^{2}/\Delta,d_{\operatorname{dim}},\Delta)$ is a homeomorphism.
To prove 3.4, we show the balls defined by $d_{\operatorname{rank}},d_{\operatorname{dim}}$ can be nested within each other. (see figure 6 for a pictorial proof)
Proof.
We’ll show the balls can be nested within one another (see figure 6), and it suffices to show this for small radius. Let $x=(x_{1},x_{2})\in\mathbb{R}_{\leq}^{2}$ and $l=x_{2}-x_{1}$. Let $0<r<\min(d_{\operatorname{dim}}(x,\Delta),d_{\operatorname{rank}}(x,\Delta))$. Note $d_{\operatorname{dim}}(x,\Delta)=l,d_{\operatorname{rank}}(x,\Delta)=\frac{1}{2}l^{2}$.
Since $r<\min(d_{\operatorname{dim}}(x,\Delta),d_{\operatorname{rank}}(x,\Delta))$, we know the shape/convexity of the balls.
Using equation 3.5, we can determine the 4 boundary points of $B_{\operatorname{rank}}(x,r)$. Let $s=\sqrt{l^{2}+2r}-l,t=l-\sqrt{l^{2}-2r}$. The boundary points are:
$(x_{1},x_{2}+s),(x_{1}+t,x_{2}),(x_{1},x_{2}-t),(x_{1}-s,x_{2})$
As shown in figure 6, when $s\geq r^{\prime}$, the 4 boundary points of $B_{r^{\prime}}(x;d_{\operatorname{dim}})$ lie inside $B_{r}(x;d_{\operatorname{rank}})$. When $t\leq r^{\prime}$, the 4 boundary points of $B_{r}(x;d_{\operatorname{rank}})$ lie inside $B_{r^{\prime}}(x;d_{\operatorname{dim}})$. To show the balls are nested, we apply implicit differentiation to equation 3.4,
$\frac{dy_{2}}{dy_{1}}=\frac{y_{2}-y_{1}}{2x_{1}-y_{1}-y_{2}}$
At $(x_{1},x_{2}-t),\frac{dy_{2}}{dy_{1}}=-1$. We apply implicit differentiation to equation 3.3,
$\frac{dy_{2}}{dy_{1}}=\frac{-2x_{2}+y_{1}+y_{2}}{y_{2}-y_{1}}$
At $(x_{1}+t,x_{2}),\frac{dy_{2}}{dy_{1}}=-1$. This can also be concluded by symmetry. As illustrated in 6,
$B_{\dim}(x,r^{\prime\prime})\subset B_{\operatorname{rank}}(x,r)\text{ if }r^{\prime\prime}\leq s=\sqrt{l^{2}+2r}-l$
and $B_{\operatorname{rank}}(x,r)\subset B_{\dim}(x,r^{\prime})$ if $t\leq r^{\prime}$. Solving for $r^{\prime}$,
$B_{\operatorname{rank}}(x,r)\subset B_{\dim}(x,r^{\prime})\text{ if }r^{\prime}\leq\min(lr-\frac{1}{2}r^{2},r).$
Figure 6: When $r<\min(d_{\operatorname{dim}}(x,\Delta),d_{\operatorname{rank}}(x,\Delta))$, the balls are of this shape. We show any ball is nested inside a ball defined by the other metric.
Proposition 3.5.
$Id:(\mathbb{R}_{\leq}^{2}/\Delta,d_{\operatorname{rank}},\Delta)\to(\mathbb{R}_{\leq}^{2}/\Delta,d_{\operatorname{dim}},\Delta)$ is not Lipschitz. Neither is $Id^{-1}$.
Proof.
It suffices to show $\frac{d_{\operatorname{rank}}(x,y)}{d_{\operatorname{dim}}(x,y)}$ and $\frac{d_{\operatorname{dim}}(x,y)}{d_{\operatorname{rank}}(x,y)}$ are unbounded. Let $x=(0,c),y\in\Delta$ for some $c>0$. Then $d_{\operatorname{rank}}(x,y)=\frac{1}{2}c^{2},d_{\operatorname{dim}}(x,y)=c$. As $c$ approaches $\infty$, $\frac{d_{\operatorname{rank}}(x,y)}{d_{\operatorname{dim}}(x,y)}=\frac{c}{2}\to\infty$. As $c$ approaches 0, $\frac{d_{\operatorname{dim}}(x,y)}{d_{\operatorname{rank}}(x,y)}=\frac{2}{c}\to\infty$. ∎
We can equip the space of finite PDs with $W_{p}^{d}$. If the 2 ground metrics $d,d^{\prime}$ are topologically equivalent, then the Wasserstein distances are also topologically equivalent.
4 Wasserstein stability from filtered chain complexes to barcodes
Let $K$ be a finite CW complex. Consider a weight $w:(K,\subset)\to([a,b],\leq)$ . For each $i\in\mathbb{N}$, we will define the $i^{th}$ chain group of $(K,w)$ as a persistence module in the Abelian category $\mathsf{Vect}^{[a,2b-a]}$. Notice we define the range of the persistence parameter to be twice the range of $w$. We will see that this interval gives the strongest stability result (proposition ).
In order to establish stability, we need a notion of distance between filtered chain complexes. Instead of taking the difference between filter functions, we observe that filtered chain complexes give rise to PMs in the following way.
Definition 4.1.
Let $K^{w}_{\bullet}$ be the sublevel filtration on $K$ with respect to $w$. Given $i\geq 0$ and a field $\mathbb{k}$, the persistence $i^{th}$ chain group of $K$ associated to $w$ is given by $C_{i}(K^{w}_{\bullet};\mathbb{k})$ in the following diagram:
$a,2b-a]\xrightarrow{K^{w}_{\bullet}}\operatorname{CW}\;\xrightarrow{C_{i}(-;\mathbb{k})}\;\operatorname{Vect_{\mathbb{k}}}$
Note this is a persistence module in $\mathsf{Vect}^{[a,2b-a]}$ where the linear maps between chain groups are induced by inclusion maps on the subcomplexes of $K$.
Since $K$ is a finite CW complex, the persistence $i^{th}$ chain group is p.f.d. and thus interval decomposable by the Structure Theorem in TDA. In particular, $C_{i}(K^{w}_{\bullet})$ is a free module generated by the $i$-simplices of $K$.
$C_{i}(K^{w}_{\bullet})=\underset{\sigma\in K_{i}}{\oplus}I_{[w(\sigma),2b-a]}$
Note each interval module in the decomposition is a free module in $\mathsf{Vect}^{[a,2b-a]}$ rather than $\mathsf{Vect}^{[a,b]}$. We denote $I_{[w(\sigma),2b-a]}$ by $I_{w(\sigma)}$, the free module generated by $\sigma$ with respect to $w$.
Remark 4.2.
If $\sigma,\tau$ are the birth and death simplices of a homology class, then the point $(w(\sigma),w(\tau))$ in the PD corresponds to the interval module $I_{w(\sigma)}/I_{w(\tau)}$.
We aim to prove the stability of the homology functor in all degrees, in other words the second map of the following diagram, from filtered chain complexes to graded persistence modules, and the grading comes from the degrees in homology.
$\operatorname{CW}^{\mathbb{R}}\overset{C_{*}(-;\mathbb{k})}{\longrightarrow}\operatorname{Ch}_{\mathbb{k}}^{\mathbb{R}}\overset{H_{*}(-)}{\longrightarrow}\underset{i\in\mathbb{N}}{\bigoplus}\mathsf{vect}^{\mathbb{R}}$
We will build up the stability theorem by first showing local stability and then extend to global stability. Below is the precise definition on what local means in this paper.
Figure 7: We plot the rank function of the persistence chain group generated by $w:K\to[0,3]$. All simplices are born in $[0,3]$ and live forever, meaning until 6.
Definition 4.3.
Let $w,v:(K,\subset)\to([a,b],\leq)$. Define partial orderings $\leq_{w},\leq_{v}$ on $K$ to be the pull back of the total ordering on $[a,b]$ w.r.t. $w,v$. We say $w,v$ are compatible if $\forall\sigma,\tau\in K,\sigma\leq_{w}\tau\Leftrightarrow\sigma\leq_{v}\tau$.
Any weight function $w:(K,\subset)\to([a,b],\leq)$ can be viewed as a point or a vector in the cube/positive cone $a,b]^{|K|}$ sitting inside the vector space $\mathbb{R}^{|K|}$. For any $\sigma\in K$, let $\pi_{\sigma}$ denote the projection of $f\in\mathbb{R}^{|K|}$ to $\mathbb{R}$, which evaluates $f$ at $\sigma$. The space $\mathbb{R}^{|K|}$ can be partitioned into chambers by the set of hyperplanes $H=\{\pi_{\sigma}x=\pi_{\tau}x|\;x\in\mathbb{R}^{|K|}\;\&\;\sigma,\tau\in K\;\&\;\sigma\neq\tau\}$. Let $S$ be a subspace of $\mathbb{R}^{|K|}$, then $S\cap H$ partitions $S$.
Lemma 4.4.
$w,v$ are compatible if and only if $w,v$ are contained in a chamber defined by $H$.
Note that the chambers are closed, so $w$ or $v$ can be on a hyperplane in $H$. If $w,v$ are compatible, then their corresponding PDs have the same set of critical simplices, with the birth and death times are determined by $w,v$.
Given $w:K\to\mathbb{R}$ and $i\in\mathbb{N}$, let $\operatorname{Dgm}_{i}(w)=H_{i}(K^{w}_{\bullet})$ and $\operatorname{Dgm}(w)=\underset{i}{\bigoplus}\operatorname{Dgm}_{i}(w)$. We compare $\operatorname{Dgm}(w),\operatorname{Dgm}(v)$ by adding up their difference at each level of the grading.
$W_{1}^{\operatorname{rank}}(\operatorname{Dgm}(w),\operatorname{Dgm}(v)):=\sum_{i}W_{1}^{\operatorname{rank}}(\operatorname{Dgm}_{i}(w),\operatorname{Dgm}_{i}(v))$
Proposition 4.5.
Let $w,v:K\to[a,b]$ be weights. If $w,v$ are compatible, then
$W_{1}^{\operatorname{rank}}(\operatorname{Dgm}(w),\operatorname{Dgm}(v))\leq\underset{\sigma\in K}{\sum}d_{\operatorname{rank}}(I_{w(\sigma)},I_{v(\sigma)})$ (4.1)
and
$W_{1}^{\operatorname{rank}}(\operatorname{Dgm}_{i}(w),\operatorname{Dgm}_{i}(v))\leq\underset{\sigma\text{ birth or death simplex}}{\sum}d_{\operatorname{rank}}(I_{w(\sigma)},I_{v(\sigma)})$ (4.2)
Proof.
Since $w,v$ are compatible, there exists a total ordering on $K$ respecting $w,v$. And the points in $\operatorname{Dgm}_{i}(w)$, $\operatorname{Dgm}_{i}(v)$ have the same set of birth and death simplex pairs, denoted by $S_{i}=\{(\sigma,\tau):\sigma,\tau\text{ form a birth-death pair}\}$. Consider a matching $\gamma$ that matches the points in Dgm($w$), Dgm($v$) according to the critical simplices.
$\begin{split}W_{1}^{\operatorname{rank}}(\operatorname{Dgm}_{i}(w),\operatorname{Dgm}_{i}(v))&\leq\operatorname{Cost}(\gamma)\\ &=\underset{(\sigma,\tau)\in S_{i}}{\sum}d_{\operatorname{rank}}((w(\sigma),w(\tau)),(v(\sigma),v(\tau)))\\ &=\underset{(\sigma,\tau)\in S_{i}}{\sum}m(T\Delta T^{\prime})\\ &\overset{(1)}{\leq}\underset{(\sigma,\tau)\in S_{i}}{\sum}m(\operatorname{supp}\>|\operatorname{rank}I_{w(\sigma)}-\operatorname{rank}I_{v(\sigma)}|\;\cup\operatorname{supp}\>|\operatorname{rank}I_{w(\tau)}-\operatorname{rank}I_{v(\tau)}|)\\ &\overset{(2)}{\leq}\underset{(\sigma,\tau)\in S_{i}}{\sum}m(\operatorname{supp}\>|\operatorname{rank}I_{w(\sigma)}-\operatorname{rank}I_{v(\sigma)}|)+m(\operatorname{supp}\>|\operatorname{rank}I_{w(\tau)}-\operatorname{rank}I_{v(\tau)}|)\\ &=\underset{(\sigma,\tau)\in S_{i}}{\sum}d_{\operatorname{rank}}(I_{w(\sigma)},I_{v(\sigma)})+d_{\operatorname{rank}}(I_{w(\tau)},I_{v(\tau)})\\ &=\underset{\sigma\text{ birth or death simplex}}{\sum}d_{\operatorname{rank}}(I_{w(\sigma)},I_{v(\sigma)})\end{split}$
(1): See figure 8
(2): The Lebesgue measure $m$ is subadditive.
Since every simplex is either a birth simplex or a death simplex of some degree of homology, summing the $i$-critical simplices over $i$ gives all simplices in $K$. Thus 4.2 implies 4.1. ∎
Figure 8: On the left we visualize the rank functions of the free modules generated by the birth simplex $\sigma$ w.r.t. $w,v$. On th right we compare $d_{\operatorname{rank}}$ to the $L^{1}$ differnce between rank functions of the free modules generated by the simplices.
Proposition 4.6.
If $w,v$ are compatible, then
$W_{1}^{\operatorname{rank}}(C_{i}(K^{w}_{\bullet}),C_{i}(K^{v}_{\bullet}))=\underset{\sigma\in K_{i}}{\sum}d_{\operatorname{rank}}(I_{w(\sigma)},I_{v(\sigma)})=\|\operatorname{rank}C_{i}(K^{w}_{\bullet})-\operatorname{rank}C_{i}(K^{v}_{\bullet})\|_{1}$
Proof.
$C_{i}(K^{w}_{\bullet}),C_{i}(K^{v}_{\bullet})$ are finitely generated free persistence modules. The interval modules admit a total order according to the birth times. By 1-dimensional optimal transport lemma, $W_{1}^{\operatorname{rank}}(C_{i}(K^{w}_{\bullet}),C_{i}(K^{v}_{\bullet}))$ is obtained by correspondences that match the order of the birth times. ∎
Corollary 4.7.
If $w,v$ are compatible, then
$W_{1}^{\operatorname{rank}}(\operatorname{Dgm}_{i}(w),\operatorname{Dgm}_{i}(v))\leq\|\operatorname{rank}C_{i}(K^{w}_{\bullet})-\operatorname{rank}C_{i}(K^{v}_{\bullet})\|_{1}+\|\operatorname{rank}C_{i+1}(K^{w}_{\bullet})-\operatorname{rank}C_{i+1}(K^{v}_{\bullet})\|_{1}$ (4.3)
and
$\overset{N}{\underset{i=n}{\sum}}W_{1}^{\operatorname{rank}}(\operatorname{Dgm}_{i}(w),\operatorname{Dgm}_{i}(v))\leq\overset{N+1}{\underset{i=n}{\sum}}\|\operatorname{rank}C_{i}(K^{w}_{\bullet})-\operatorname{rank}C_{i}(K^{v}_{\bullet})\|_{1}$
Theorem 4.8.
Let $w,v:K\to[a,b]$ be weights. Let $\operatorname{Dgm}(w)=\underset{i}{\bigoplus}H_{i}(K^{w}_{\bullet})$. For each $i\in\mathbb{N}$, let $\operatorname{Dgm}_{i}(w)=H_{i}(K^{w}_{\bullet})$. Let $\operatorname{Dgm}(v)$ and $\operatorname{Dgm}_{i}(v)$ be defined similarly.
$W_{1}^{\operatorname{rank}}(\operatorname{Dgm}(w),\operatorname{Dgm}(v))\leq\underset{\sigma\in K}{\sum}d_{\operatorname{rank}}(I_{w(\sigma)},I_{v(\sigma)})$ (4.4)
and
$W_{1}^{\operatorname{rank}}(\operatorname{Dgm}_{i}(w),\operatorname{Dgm}_{i}(v))\leq\underset{\sigma\in K_{i}\cup K_{i+1}}{\sum}d_{\operatorname{rank}}(I_{w(\sigma)},I_{v(\sigma)})$ (4.5)
The proof of 4.8 generalizes from the proof of 4.5. Now we look at the case where the weights $w_{0},w_{1}$ on $K$ are not compatible. Consider the convex combinations of $w_{0}$ and $w_{1}$, $w_{t}:=(1-t)w_{0}+t\;w_{1}$ for $t\in[0,1]$. Let $l=\{w_{t}|\;t\in[0,1]\}$. For each hyperplane $h\in H,l\cap h$ is either empty, one point or the entire line segment $l$. (Shown below if needed)
Lemma 4.9.
Consider a straight line $l$ and a subspace $h$ in $\mathbb{R}^{n}$ for some $n$. If $l\cap h$ is non-empty, $l\cap h$ is either one point or the entire line segment $l$.
Proof.
$l\cap h$ is convex since $l$ and $h$ are convex subsets of $\mathbb{R}^{|K|}$. Convex subsets of $l$ are either one point or a line segment $j\subseteq l$. If $l\cap h$ is a proper subset of $l$ and not a single point, then $l\setminus j$ is contained in $h$ by closure under addition since $h$ is a subspace. ∎
In particular, $l$ crosses any hyperplane at most once. Since $K$ is a finite CW complex, $|H|<\infty$. Then $Q:=\{l\cap h|\;h\in H\;\&\;l\not\subset h\}$ consists of finitely many points. Define $P=Q\cup\{w_{0},w_{1}\}$. Order $P$ into $\{a_{j}\}_{j=0}^{N}$ based on the values of $t$ and let $a_{0}=w_{0},a_{N}=w_{1}$. Now $P$ is a partition of $l$. For notational simplicity, we denote the line segment from $a_{j}$ to $a_{k}$ by $a_{j},a_{k}]$ for $0\leq j,k\leq N$.
Observe that for each $0\leq j<N$, all weights $w_{t}\in[a_{j},a_{j+1}]$ are compatible with each other.
Proof.
Let $w_{0}=w,w_{1}=v$.
[…]
(1) Triangle inequality
(2) Birth simplices are in $K_{i}$ and death simplices are in $K_{i+1}$
(3) For any $\sigma\in K$, $(a_{j}(\sigma))_{j=0}^{N}$ is a monotone sequence. Recall $T_{I}$ is the shaded triangle in fig 1. If $a_{0}(\sigma)\geq a_{N}(\sigma)$, $\forall j\in[0,N-1],d_{\operatorname{rank}}(I_{a_{j}(\sigma)},I_{a_{j+1}(\sigma)})=m(T_{I_{a_{j}(\sigma)}})-m(T_{I_{a_{j+1}(\sigma)}})$, and vice versa for $a_{0}(\sigma)\leq a_{N}(\sigma)$.
Similar to proposition 4.5, equation 4.5 implies $4.$ ∎
Lemma 4.10.
Given $w,v:(K,\subset)\to([a,b],\leq)$. Let $C_{*}(K^{w}_{\bullet}),C_{*}(K^{v}_{\bullet})\in\mathsf{vect}^{[a,c]}$ be the persistence chain groups. 4.8 holds if $c\geq 2b-a$.
Proof.
Consider 2 PDs containing $(a,\frac{a+b}{2})$ and $(a,b)$ respectively. From figure 8 we can see the stability result holds only if $c\geq 2b-a$. ∎
Definition 4.11.
For $p\in[1,\infty]$, define $d_{\operatorname{rank},p}(I,J)=(\int_{\mathbb{R}^{2}}|\operatorname{rank}\>I-\operatorname{rank}\>J|^{p})^{1/p}$. Denote the $p$-Wasserstein distance with $d_{\operatorname{rank},p}$ by $W_{p}^{\operatorname{rank},p}$.
Let $g_{p}:\sigma\mapsto d_{\operatorname{rank},p}(I_{w(\sigma)},I_{v(\sigma)})$. We can rewrite 4.8 as:
$W_{1}^{\operatorname{rank}}(\operatorname{Dgm}(w),\operatorname{Dgm}(v))\leq\|g_{1}\|_{1}$ (4.6)
Furthermore, we can generalize the stability result 4.8 to $p$-Wasserstein distances with appropriate metrics.
Theorem 4.12.
For $p\in[1,\infty]$,
$W_{p}^{\operatorname{rank},p}\operatorname{Dgm}_{i}(w),\operatorname{Dgm}_{i}(v))\leq\|g_{p}|_{K_{i}\cup K_{i+1}}\|_{p}$ (4.7)
and
$\underset{i\geq 0}{\sum}W_{p}^{\operatorname{rank},p}\operatorname{Dgm}_{i}(w),\operatorname{Dgm}_{i}(v))\leq\|g_{p}\|_{p}$ (4.8)
Proposition 4.13.
Given $w,v:K\to[a,b]$. The map from $CW^{[c,d]}$ to $\mathsf{vect}_{\mathbb{k}}$ is stable w.r.t. $W_{1}^{\operatorname{rank}}$ iff $c\leq a$ and $d\geq 2b-a$.
4.1 Comparison with previous stability results
Let $a,b\in\mathbb{R}$ and $a<b$. Let $IM$ denote the set of interval modules in $\mathsf{vect}^{[a,b]}$. For all $x\in[a,b]$, let $I_{x}$ be the free module with support $[x,b]$. $d_{\operatorname{rank}}$ and $d_{\operatorname{dim}}$ are metrics on $IM$. Note that $d_{\operatorname{dim}}(I_{x},I_{x^{\prime}})=|x-x^{\prime}|$, independent of $b$, whereas $d_{\operatorname{rank}}(I_{x},I_{x^{\prime}})=|\frac{1}{2}(b-x)^{2}-\frac{1}{2}(b-x^{\prime})^{2}|$ depends on $b$.
We rewrite the metric induced by the $p$-norm on $\mathbb{R}^{2}$ using $d_{\operatorname{dim}}$ in the following way.
Consider interval modules $I,I^{\prime}$ with supports $x,y],[x^{\prime},y^{\prime}]$ respectively. For $p\in[1,\infty]$, let $d_{p}(I,I^{\prime})=\|(d_{\operatorname{dim}}(I_{x},I_{x^{\prime}}),d_{\operatorname{dim}}(I_{y},I_{y^{\prime}}))\|_{p}$ . Observe that $d_{p}(I,I^{\prime})=\|(x,y)-(x^{\prime},y^{\prime})\|_{p}$.
Let $f:\sigma\mapsto d_{\operatorname{dim}}(I_{w(\sigma)},I_{v(\sigma)})$. Observe f=wvf=w-v. We can rewrite the bottleneck stability ([12]) as:
$W_{\infty}^{d_{\infty}}(\operatorname{Dgm}_{i}(w),\operatorname{Dgm}_{i}(v))\leq\|f\|_{\infty}$ (4.9)
for some $i\in\mathbb{N}$, and $d_{\infty}(I_{[b,d]},I_{[b^{\prime},d^{\prime}]})$ can be expressed as
$\|(d_{\operatorname{dim}}(I_{b},I_{b^{\prime}}),d_{\operatorname{dim}}(I_{d},I_{d^{\prime}}))\|_{\infty}$ (4.10)
The $p$-Wasserstein stability of barcodes shown in [33] can be reformulated as:
$W_{p}^{d_{p}}(\operatorname{Dgm}(w),\operatorname{Dgm}(v))\leq\|f\|_{p}$ (4.11)
Let $g:\sigma\mapsto d_{\operatorname{rank}}(I_{w(\sigma)},I_{v(\sigma)})$. We rewrite 4.8 as:
$W_{1}^{\operatorname{rank}}(\operatorname{Dgm}(w),\operatorname{Dgm}(v))\leq\|g\|_{1}$ (4.12)
5 Stability of vectorization using persistence landscape
In practice, one often needs to conduct analysis on PDs. One common way is to map PDs into vector spaces for applying machine learning tools. Bubenik’s Persistence Landscapes are widely used for this purpose. PLs are non-negative functions on $\mathbb{N}\times\mathbb{R}$. We show stability of PLs using our new metric. Furthermore, the bound is tight.
Definition 5.1.
Let $\alpha$ be a formal sum of points in $\mathbb{R}_{\leq}^{2}$. Define the persistence landscape (PL) of $\alpha$ to be a function from $\mathbb{N}\times\mathbb{R}$ to $0,\infty)$ where $\lambda(a,t)=sup\{h>0:|\{(x_{1},x_{2})\in\alpha:x_{1}\leq t-h,x_{2}\geq t+h\}|\geq a\}$.Equivalently, $\lambda(a,t)=sup\{h>0:\operatorname{rank}\>M(t-h\leq t+h)\geq a\}$
Let $\Lambda$ denote the vectorization map that sends PDs to PLs.
5.1 $L^{1}$-stability for persistence landscapes
Bubenik has shown vectorization using PLs is injective, invertible, and $L^{\infty}$-stable with respect to the bottleneck distance([3]). We’ll show the vectorization is also stable with respect to the 1-Wasserstein distance with the metric $d_{\operatorname{rank}}$.
Theorem 5.2.
Let $\alpha,\beta$ be (finite) formal sums on $\mathbb{R}_{\leq}^{2}$. Then $\|\Lambda\alpha-\Lambda\beta\|_{1}\leq\frac{1}{2}W_{1}^{\operatorname{rank}}(\alpha,\beta)$.
Proof.
Since the PDs have finitely many points, we can find an optimal coupling $\gamma=\sum_{k=1}^{N}(x_{k},y_{k})$.
$\begin{split}\|\Lambda\alpha-\Lambda\beta\|_{1}&=\int_{\mathbb{N}\times\mathbb{R}}|\lambda(k,t)-\lambda^{\prime}(k,t)|d(k,t)\\ &=\sum_{k=1}^{N}\int_{\mathbb{R}}|\lambda(k,t)-\lambda^{\prime}(k,t)|dt\\ &=\int_{\mathbb{R}}\sum_{k=1}^{N}|\lambda(k,t)-\lambda^{\prime}(k,t)|dt\\ &=\int_{\mathbb{R}}\sum_{k=1}^{N}|u_{(k)}(t)-v_{(k)}(t)|dt\\ &\overset{(1)}{\leq}\int_{\mathbb{R}}\sum_{k=1}^{N}|u_{k}(t)-v_{k}(t)|dt\\ &=\int_{\mathbb{R}}\sum_{k=1}^{N}|\Lambda\{(x_{k})\}(t)-\Lambda\{(y_{k})\}(t)|dt\\ &=\sum_{k=1}^{N}\int_{\mathbb{R}}|\Lambda\{(x_{k})\}(t)-\Lambda\{(y_{k})\}(t)|dt\\ &=\sum_{k=1}^{N}d_{\operatorname{rank}}(x_{k},y_{k})\\ &=W_{1}^{\operatorname{rank}}(alpha,\beta)\end{split}$
(1): $\forall t\in\mathbb{R}$, define $u_{k}(t):=\Lambda\{x_{k}\}(1,t),v_{k}(t):=\Lambda\{y_{k}\}(1,t)$. And for each $t$, permute $u_{k}(t)$ and order $u_{\{}(k)\}(t)$ such that $u_{\{}(1)\}(t)\geq u_{\{}(2)\}(t)\geq u_{\{}(3)\}(t)\geq...\geq u_{\{}(N)\}(t)$. Similarly obtain $v_{\{}(k)\}(t)$ for $1\leq k\leq N$. Note $\lambda(k,t)=u_{(k)}(t),\lambda^{\prime}(k,t)=v_{(k)}(t)$.
The inequality comes from a lemma in 1-dimensional optimal transport.
5.2 Landscape stability for countable persistence diagrams
We’ll gather some facts and propositions in order to extend the result to countably supported PDs.
Lemma 5.3.
Let $d$ be a metric on interval modules. Let $\alpha,\beta,\gamma\in\bar{D}_{1}(\mathbb{R}_{\leq}^{2},\Delta)$ be countable PDs with finite 1-Wasserstein distances. Then $W_{1}^{d}(\alpha,\gamma)\leq W_{1}^{d}(\alpha,\beta)+W_{1}^{d}(\beta,\gamma)$.
Lemma 5.4.
If $\{\alpha_{n}\}_{n=1}^{\infty}$ is an increasing sequence of formal sums. Then $\{W_{1}^{d}(\alpha_{n},0)\}_{n=1}^{\infty}$ is an increasing sequence of nonnegative real numbers. If $\alpha=\underset{n\to\infty}{lim}\>\alpha_{n}$, and $W_{1}^{d}(\alpha,0)<\infty$, then $\{W_{1}^{d}(\alpha_{n},0)\}_{n=1}^{\infty}$ is bounded above and $W_{1}^{d}(\alpha_{n},0)\to W_{1}^{d}(\alpha,0)$.
Proposition 5.5.
If $W_{1}^{d}(\alpha,0)<\infty$, then $\|\Lambda\alpha\|_{1}=\frac{1}{2}W_{1}^{\operatorname{rank}}(\alpha,0)$.
Proposition 5.6.
Assume $W_{1}^{d}(\alpha,0)<\infty$, $W_{1}^{d}(\beta,0)<\infty$, $W_{1}^{d}(\gamma,0)<\infty$. Then $W_{1}^{d}(\alpha,\beta)\leq W_{1}^{d}(\alpha,\gamma)+W_{1}^{d}(\gamma,\beta)$.
Lemma 5.7.
Assume $\alpha\geq\beta$ as countable formal sums. Then $W_{1}^{d}(\alpha,\beta)=W_{1}^{d}(\alpha-\beta,0)=W_{1}^{d}(\alpha,0)-W_{1}^{d}(\beta,0)$.
Lemma 5.8.
Assume $\alpha\geq\beta$ as countable formal sums. Then $\|\Lambda\alpha-\Lambda\beta\|_{1}=\|\Lambda\alpha\|_{1}-\|\Lambda\beta\|_{1}$.
Now we are ready to extend the stability result to countable persistence diagrams.
Theorem 5.9.
Let $\alpha,\beta$ be 1-finite countable formal sums on $\mathbb{R}_{\leq}^{2}$. Then $\|\Lambda\alpha-\Lambda\beta\|_{1}\leq\frac{1}{2}W_{1}^{\operatorname{rank}}(\alpha,\beta)$.
Proof.
Let $\alpha=\underset{i\geq 1}{x_{i}},\beta=\underset{i\geq 1}{y_{i}}$, and for all $k\in\mathbb{N}$, $\alpha_{k}=\underset{1\leq i\leq k}{x_{i}},\beta_{k}=\underset{1\leq i\leq k}{y_{i}}$.
Let $\epsilon>0$. We’ll show $\|\Lambda\alpha-\Lambda\beta\|_{1}<\frac{1}{2}W_{1}^{\operatorname{rank}}(\alpha,\beta)+\epsilon$,
By 5.4, we can fix NN such that $W_{1}^{\operatorname{rank}}(\alpha,\alpha_{N})<\epsilon,W_{1}^{\operatorname{rank}}(\beta,\beta_{N})<\epsilon$.
[…]
6 Graded 1-Wasserstein distances
We recall the construction of graded PDs from [2] in appendix. In short, PDs can be obtained by taking the Möbius inversion of the rank functions, and graded PDs are obtained by taking the Möbius inversion of the graded rank functions. And the grading is on the value of the rank function, with integral value.
6.1 Useful results
Lemma 6.1.
Let $\alpha,\beta,\sigma,\tau$ be finite PDs. Let $d$ be a metric. Then
$W_{1}^{d}(\alpha+\sigma,\beta+\tau)\leq W_{1}^{d}(\alpha,\beta)+W_{1}^{d}(\sigma,\tau)$ (6.1)
and
$W_{1}^{d}(\alpha+\sigma,\beta+\sigma)=W_{1}^{d}(\alpha,\beta)$ (6.2)
Corollary 6.2.
Let $\alpha,\beta,\sigma,\tau,\gamma,\eta$ be finite PDs. Let $d$ be a metric.
$W_{1}^{d}(\alpha+\gamma,\beta+\sigma)+W_{1}^{d}(\tau+\sigma,\eta+\gamma)\geq W_{1}^{d}(\alpha+\tau,\beta+\eta)$
Proof.
The proof follows from 6.1 and 6.2. ∎
6.2 Stability of graded PDs
Each PD $\alpha$ has a sequence of functions $\alpha^{(k)}:\mathbb{R}_{\leq}^{2}\to\{-1,0,1\}$, called graded PDs. Let $\alpha_{+}^{(k)},\alpha_{-}^{(k)}$ be the positive and negative parts of the graded PDs.
We show another version of the stability theorem in terms of the 1-Wasserstein distance for graded PDs.
We show that the graded Wasserstein distances are more discriminative than the Wasserstein distance in the following sense.
Proposition 6.3.
Let $\alpha,\beta$ be PDs of p.f.d. persistence modules. Let $d$ be a metric on interval modules. Then, $\exists N\in\mathbb{Z}$ such that
$W_{1}^{d}(\alpha,\beta)\leq\underset{1\leq k\leq N}{\sum}W_{1}^{d}(\alpha^{(k)},\beta^{(k)})$
Corollary 6.4.
$\|\Lambda\alpha-\Lambda\beta\|_{1}\leq\underset{1\leq k\leq N}{\sum}W_{1}^{\operatorname{rank}}(\alpha^{(k)},\beta^{(k)})$
Appendix A Appendix
A.1 Triangle inequality for Wasserstein distances on countable PDs
Proof.
[…]
$L\subset I,\phi:L\hookrightarrow J,M\subset J,\psi:M\hookrightarrow K$
and $\hat{\tau}=\underset{l\in L}{\sum}(x_{l},y_{\phi(l)})+\underset{i\in I\setminus L}{\sum}(x_{i},a_{i})+\underset{j\in J\setminus\phi(L)}{\sum}(b_{j},y_{j})$,
$\hat{\sigma}=\underset{m\in M}{\sum}(y_{m},z_{\phi(m)})+\underset{j\in J\setminus M}{\sum}(y_{j},c_{j})+\underset{k\in K\setminus\psi(M)}{\sum}(d_{k},z_{k})$
where $\{a_{i}\in\Delta:i\in I\setminus L,d(x_{i},a_{i})=d(x_{i},\Delta)\},\{b_{j}:j\in J\setminus\phi(L),d(y_{j},b_{j})=d(y_{j},\Delta)\},\{c_{j}:j\in J\setminus M,d(y_{j},c_{j})=d(y_{j},\Delta)\},\{d_{k}:k\in K\setminus\psi(M),d(z_{k},d_{k})=d(z_{k},\Delta)\}$.
Figure 9: construct a coupling $\zeta$ by composing couplings $\phi,\psi$
Define $\hat{\pi}$ in the following way as in figure 9: Observe $\phi(L)\cap M\subset J$. Let $N=\phi^{-1}(\phi(L)\cap M)=L\cap\phi^{-1}(M)\subset I$. Let $\zeta:N\hookrightarrow K$ where $\zeta(n)=\psi(\phi(n))$. Define $\hat{\pi}=\underset{n\in N}{\sum}(x_{n},z_{\zeta(m)})+\underset{i\in I\setminus N}{\sum}(x_{i},e_{i})+\underset{k\in K\setminus\zeta(N)}{\sum}(f_{k},z_{k})$ where $\{e_{i}:i\in I\setminus N\},\{f_{k}:k\in K\setminus\zeta(N)\}\subset\Delta$.
[…]
$\overset{(1)}{\leq}$ $\underset{n\in N}{\sum}(d(x_{n},y_{\phi(n)})+d(y_{\phi(n)},z_{\zeta(m)}))\;+$ $\underset{i\in I\setminus L}{\sum}d(x_{i},e_{i})$ $+\underset{i\in L\setminus N}{\sum}d(x_{i},e_{i})+$ $\underset{k\in K\setminus\psi(M)}{\sum}d(f_{k},z_{k})$ $+\underset{k\in\psi(M)\setminus\zeta(N)}{\sum}d(f_{k},z_{k})$
(1): we expand the first term by triangle inequality of $d$, the second term by $I\setminus N=(I\setminus L)\sqcup(L\setminus N)$, and the third term by $K\setminus\zeta(N)=(K\setminus\psi(M))\sqcup(\psi(M)\setminus\zeta(N))$.
[…]
$\begin{split}&\operatorname{Cost}(\hat{\tau})+\operatorname{Cost}(\hat{\sigma})-\operatorname{Cost}(\hat{\pi})=\\ &\underset{l\in L}{\sum}d(x_{l},y_{\phi(l)})+\underset{i\in I\setminus L}{\sum}d(x_{i},a_{i})+\underset{j\in J\setminus\phi(L)}{\sum}d(b_{j},y_{j})+\underset{m\in M}{\sum}d(y_{m},z_{\phi(m)})+\underset{j\in J\setminus M}{\sum}d(y_{j},c_{j})+\underset{k\in K\setminus\psi(M)}{\sum}d(d_{k},z_{k})\\ &\;-(\underset{n\in N}{\sum}(d(x_{n},y_{\phi(n)})+d(y_{\phi(n)},z_{\zeta(m)}))+\underset{i\in I\setminus L}{\sum}d(x_{i},a_{i})+\underset{i\in L\setminus N}{\sum}d(x_{i},a_{i})+\underset{k\in K\setminus\psi(M)}{\sum}d(f_{k},z_{k})+\underset{k\in\psi(M)\setminus\zeta(N)}{\sum}d(f_{k},z_{k}))\\ \\ =&\underset{l\in L\setminus N}{\sum}(x_{l},y_{\phi(l)})+\underset{j\in J\setminus\phi(L)}{\sum}(b_{j},y_{j})+\underset{m\in M\setminus\phi(N)}{\sum}d(y_{m},z_{\phi(m)})+\underset{j\in J\setminus M}{\sum}d(y_{j},c_{j}))-(\underset{i\in L\setminus N}{\sum}d(x_{i},e_{i})+\underset{k\in\psi(M)\setminus\zeta(N)}{\sum}d(f_{k},z_{k}))\end{split}$
Observe that by construction of $N$, $\underset{n\in N}{\sum}d(x_{n},y_{\phi(n)})\leq\underset{n\in L}{\sum}d(x_{n},y_{\phi(n)})$ since $N\subset L$, and $\underset{n\in N}{\sum}d(y_{\phi(n)},z_{\zeta(m)})\leq\underset{m\in M}{\sum}d(y_{m},z_{\phi(m)})$ since $\phi(N)\subset M$.
And $\underset{i\in I\setminus L}{\sum}d(x_{i},\Delta)+\underset{k\in K\setminus\psi(M)}{\sum}d(z_{k},\Delta)$ cancel out in the difference.
To show the expression is non-negative, we observe that $\forall l\in L\setminus N,d(x_{l},\Delta)\leq d(x_{l},y_{\phi(l)})+d(y_{\phi(l)},\Delta)$, since there exists a unique $j\in J$ such that $j=\phi(l)$, if $L\setminus N\neq\emptyset$. Similarly, $\forall k\in\psi(M)\setminus\zeta(N),d(z_{k},\Delta)\leq d(y_{m},z_{\psi(m)})+d(y_{m},\Delta)$, since there exists a unique $m\in M\setminus\phi(N)$ such that $\psi(m)=k$ and $m\notin\phi(L)$, if $\psi(M)\setminus\zeta(N)\neq\emptyset$.
A.2 Construction of graded PDs
Definition A.1.
Let $(P,\leq)$ be a locally finite partially ordered set. The Möbius function $\mu$ of $P$ is defined recursively as follows:
  1. 1.
    $\forall x\in P,\mu(x,x)=1$
  2. 2.
    $\forall x,y\in P,x<y\implies\underset{x\leq t\leq y}{\sum}\mu(x,t)=0$
Example A.2.
Consider the totally ordered set $\mathbb{Z}$ with $\leq$ .
[…]
Example A.3.
Consider the partially ordered set $\operatorname{Int}(\mathbb{Z})$ with $\subset$ .
[…]
Example A.4.
Consider the partially ordered set $\mathbb{Z}\times\mathbb{Z}$ with the product order. Notice $\operatorname{Int}(\mathbb{Z})=\mathbb{Z}^{\text{op}}\times\mathbb{Z}$ in the previous example.
[…]
Let $M$ be a p.f.d. module. For any $r,s]\in\operatorname{Int}(\mathbb{R})$, $\operatorname{rank}M_{r\leq s}$ denotes the rank of the linear map $M_{r\leq s}$. The rank invariant of $M$ is an integral function on $\operatorname{Int}(\mathbb{R})$ given below:
$\begin{split}\operatorname{rank}M:\operatorname{Int}(\mathbb{R})&\to\mathbb{Z}\\ [r,s]&\mapsto\operatorname{rank}M_{r\leq s}\end{split}$
Define the $i^{th}$-graded rank function as follows:
$\begin{split}(\operatorname{rank}M)_{i}:\operatorname{Int}([a,b])&\to\{0,1\}\\ [r,s]&\mapsto 1\text{ iff }\operatorname{rank}M_{r\leq s}\geq i\end{split}$
Fact A.5.
$\underset{i=1}{\overset{N}{\sum}}(\operatorname{rank}M)_{i}=\operatorname{rank}M$ for some $N\in\mathbb{Z}$.
Let $\mu$ be the Möbius function on the partially ordered set $(\operatorname{Int}(P),\subset)$, given in example A.3. Viewing PDs as integral functions on $\mathbb{R}_{\leq}^{2}$, the author in [29] expressed PDs as:
$\operatorname{Dgm}(F)(r,s)=\underset{[a,b]\subset[r,s]}{\sum}\mu([a,b],[r,s])\operatorname{rank}M([r,s])$
where $(r,s)\in\mathbb{R}_{\leq}^{2}$. The authors in [2] defined graded PDs by taking the Möbius inversion of the graded rank function instead.
$\operatorname{Dgm}_{i}(F)(r,s)=\underset{[a,b]\subset[r,s]}{\sum}\mu([a,b],[r,s])(\operatorname{rank}M)_{i}([r,s])$
Graded PDs are formal sums into $\{0,1,-1\}$, i.e. PDs with points of multiplicity 1 or -1. By fact A.5, $\operatorname{Dgm}(F)=\underset{i\geq 1}{\sum}\operatorname{Dgm}_{i}(F)$.
A.3 Wasserstein distance for graded PDs
Lemma A.6.
Let $\alpha,\beta,\sigma,\tau$ be finite PDs. Let $d$ be a metric. Then
$W_{1}^{d}(\alpha+\sigma,\beta+\tau)\leq W_{1}^{d}(\alpha,\beta)+W_{1}^{d}(\sigma,\tau)$ (A.1)
and
$W_{1}^{d}(\alpha+\sigma,\beta+\sigma)=W_{1}^{d}(\alpha,\beta)$ (A.2)
Proof.
Proof to A.1:
By assumption, $\alpha,\beta,\sigma,\tau$ are finite PDs, then $\exists$ optimal couplings $\gamma_{1},\gamma_{2},\gamma_{3}$, which are finite discrete matchings, such that:
$W_{1}^{d}(\alpha,\beta)=\operatorname{Cost}(\gamma_{1})=\underset{(x.y)\in\gamma_{1}}{\sum}d(x,y)$
$W_{1}^{d}(\sigma,\tau)=\operatorname{Cost}(\gamma_{2})=\underset{(x.y)\in\gamma_{2}}{\sum}d(x,y)$
$W_{1}^{d}(\alpha+\sigma,\beta+\tau)=\operatorname{Cost}(\gamma_{3})=\underset{(x.y)\in\gamma_{3}}{\sum}d(x,y)\\$
Note $\gamma_{1}+\gamma_{2}$ defines a coupling between $\alpha+\sigma$ and $\beta+\tau$. Then by construction, $\operatorname{Cost}(\gamma_{3})\leq\underset{(x.y)\in\gamma_{1}+\gamma_{2}}{\sum}d(x,y)=\operatorname{Cost}(\gamma_{1})+\operatorname{Cost}(\gamma_{2})$.
Proof to A.2:
Since $\sigma$ is a finite PD, it suffices to show the case when $\sigma$ = $x_{0}$ for some $x_{0}\in\mathbb{R}_{\leq}^{2}$. Then by induction on the number of points in $\sigma$, the result holds for finite $\sigma$.
Let $\gamma_{1},\gamma_{2}$ be optimal couplings where $W_{1}^{d}(\alpha,\beta)=\operatorname{Cost}(\gamma_{1}),W_{1}^{d}(\alpha+\sigma,\beta+\sigma)=\operatorname{Cost}(\gamma_{2})$.
There’s a canonical injection/inclusion map from the couplings between $\alpha,\beta$ to the couplings between $\alpha+\sigma,\beta+\sigma$, given by $\gamma_{1}\hookrightarrow\gamma_{1}+(x_{0},x_{0})=:\bar{\gamma_{1}}$. Observe $\operatorname{Cost}(\gamma_{1})=\operatorname{Cost}(\bar{\gamma_{1}})$ since $d(x_{0},x_{0})=0$. Since $\gamma_{2}$ is optimal, $\operatorname{Cost}(\bar{\gamma_{1}})\geq\operatorname{Cost}(\gamma_{2})$.
On the other hand, if $x_{0}$ is not matched to itself, then $\exists y_{0}\in\beta,z_{0}\in\alpha$ such that $(x_{0},y_{0}),(z_{0},x_{0})\in\gamma_{2}$.
[…]
We define $\tilde{\gamma_{2}}$ to be the coupling that replaces $(x_{0},y_{0}),(z_{0},x_{0})$ by $(z_{0},y_{0})$ in $\gamma_{2}$. But since $\gamma_{2}$ is an optimal coupling, $\operatorname{Cost}(\gamma_{2})=\operatorname{Cost}(\tilde{\gamma_{2}})+d(x_{0},x_{0})$. Note $\tilde{\gamma_{2}}$ is a coupling between $\alpha,\beta$ and because $\gamma_{1}$ is an optimal coupling, $\operatorname{Cost}(\tilde{\gamma_{2}})\geq\operatorname{Cost}(\gamma_{1}).$ Then $\operatorname{Cost}(\gamma_{2})\geq\operatorname{Cost}(\gamma_{1}).$
Corollary A.7.
Let $\alpha,\beta,\sigma,\tau,\gamma,\eta$ be finite PDs. Let $d$ be a metric.
$W_{1}^{d}(\alpha+\gamma,\beta+\sigma)+W_{1}^{d}(\tau+\sigma,\eta+\gamma)\geq W_{1}^{d}(\alpha+\tau,\beta+\eta)$
Proof.
The proof follows from A.1 and A.2. ∎
Proof.
(proof to 6.3) Note p.f.d persistence modules have bounded rank functions. Fix $N$ to be an upper bound of the rank functions. Then $\alpha=\underset{1\leq k\leq N}{\sum}\alpha^{(k)},\beta=\underset{1\leq k\leq N}{\sum}\beta^{(k)}$.
$\begin{split}\underset{1\leq k\leq N}{\sum}W_{1}^{d}(\alpha^{(k)},\beta^{(k)})&\overset{(1)}{=}\underset{1\leq k\leq N}{\sum}W_{1}^{d}(\alpha^{(k)}_{+}+\beta^{(k)}_{-},\beta^{(k)}_{+}+\alpha^{(k)}_{-})\\ &\overset{(2)}{\geq}W_{1}^{d}(\underset{1\leq k\leq N}{\sum}\alpha^{(k)},\underset{1\leq k\leq N}{\sum}\beta^{(k)})\\ &=W_{1}^{d}(\alpha,\beta)\end{split}$
(1): By definition of Wasserstein distance for signed PDs
(2): The graded PDs add up to the PD, so the points in $\alpha^{(k)}_{-},\beta^{(k)}_{-}$ will be canceled out by the positive part of the graded PDs, $\alpha^{(k)}_{+},\beta^{(k)}_{+}$. Then we can apply 6.2 to obtain the inequality. ∎